A urine sample is diluted 1:100. 200 µl of the 1:100 dilution is spread onto an agar plate. After incubation, 250 colonies are observed. What is the concentration of bacteria in the original urine sample? Show your work and include proper units for full credit.
Answer #1

CFU/mL= number of colonies/( volume plated in mL \times dilution)

dilution= 1/100

number of colonies= 250

volume plated= 200 uL

1mL= 1000 uL

200 uL= 200/1000=0.2 mL

volume plated in mL= 0.2 mL

CFU/mL= 250/(0.2\times(1/100))

= 125000 /mL

concentration of bacteria in the original urine sample = 125000 cells/mL